### v1

# Frustrated Random Walks: A Fast Method to Compute Node Distances on Hypergraphs

Abstract
A hypergraph is a generalization of a graph that arises naturally when attribute-sharing among entities is considered. Although a hypergraph can be converted into a graph by expanding its hyperedges into fully connected subgraphs, going the reverse way is computationally complex and NP-complete. We therefore hypothesize that a hypergraph contains more information than a graph. In addition, it is more convenient to manipulate a hypergraph directly, rather than expand it into a graph. An open problem in hypergraphs is how to accurately and efficiently calculate their node distances. Estimating node distances enables us to find a node’s nearest neighbors, and perform label propagation on hypergraphs using a K-nearest neighbors (KNN) approach. In this paper, we propose a novel approach based on random walks to achieve label propagation on hypergraphs. We estimate node distances as the expected hitting times of random walks. We note that simple random walks (SRW) cannot accurately describe highly complex real-world hypergraphs, which motivates us to introduce frustrated random walks (FRW) to better describe them. We further benchmark our method against DeepWalk, and show that while the latter can achieve comparable results, FRW has a distinct computational advantage in cases where the number of targets is fairly small. For such cases, we show that FRW runs in significantly shorter time than DeepWalk. Finally, we analyze the time complexity of our method, and show that for large and sparse hypergraphs, the complexity is approximately linear, rendering it superior to the DeepWalk alternative.
I Introduction
A graph is a useful data structure for describing complex relations in real world. In a graph, nodes are connected by edges, and each edge contains exactly two nodes. A hypergraph is an extension of ordinary graphs where each hyperedge of a hypergraph may contain an arbitrary number of nodes (see Fig. 5 in Appendix A for an illustration). In the special case when each hyperedge contains just two nodes, the hypergraph reduces to a graph. Each hyperedge of a hypergraph is considered a clique in a graph. We can therefore convert a hypergraph into a graph by expanding each hyperedge. However, hypergraphs contain more information than what can be encapsulated into its expanded graphs. If we simply expand a hypergraph into a graph, the explicit information of the cliques is lost. Although the conversion from hypergraphs to graphs is reversible in theory, the NP-completeness of detecting maximal cliques in a graphs renders it irreversible in practice. Furthermore, the expansion of hyperedges can be computationally expensive and memory-intensive considering that a hyperedge of $N$ nodes yields $N(N-1)/2$ undirected edges in a graph. It is hence advantageous to directly work with hypergraphs whenever possible rather than with their expanded graphs.
One of the central problems in graph theory is the quantitative determination of graph node distances. The literature has multiple algorithms that aim to achieve this, such as DeepWalkPerozzi et al. (2014), graphSAGEHamilton et al. (2017), hitting times of random walksKlein and Randić (1993) and frustrated random walksLi and Le (2020), among others. We can easily generalize DeepWalk to hypergraphs, yet for the other algorithms, such generalization is highly non-trivial. In this paper, we generalize frustrated random walks to hypergraphs, and show that the approach is on par with the performance of DeepWalk even for complex hypergraphs. There are four reasons that motivate using frustrated random walks (FRW) as opposed to DeepWalk. First, in applications where the goal is to find the nearest neighbors of a few nodes in a large hypergraph, FRW is a preferable and faster option than DeepWalk. Second, FRW conveniently gives a closed-form and interpretable solution for node distances, a solution that is impossible to obtain using any deep-learning based method. Third, FRW does not require any parameter tuning, whereas DeepWalk requires heavy investment in parameter tuning. Finally, and unlike in DeepWalk, the node distances of FRW are asymmetric, rendering it more suitable to describe real-world relationships which are generally asymmetric and non-reflective.
More formally, a hypergraph consists of nodes and hyperedges. Each hyperedge is a subset of the node set $\mathbb{V}$. For clarity purpose, we use in this paper Latin letters to indicate node indices and Greek letters to indicate hyperedge indices. The incidence matrix of a hypergraph is defined as:
$e_{i\alpha}=\begin{cases}&w,\text{if vertex }v_{i}\in E_{\alpha}\\ &0,\text{otherwise}\end{cases}$ (1)
In the above definition, $w$ is the weight of vertex $v_{i}$ in hyperedge $E_{\alpha}$. If the hypergraph is unweighted, then $w=1$ always holds. If we think of $v_{i}$ as a member of a community $E_{\alpha}$, then $e_{i\alpha}$ can be thought of as the loyalty of $v_{i}$ to $E_{\alpha}$. By definition, the degree of a node $v_{i}$ is $D_{i}=\sum_{\alpha}e_{i\alpha}$, and the degree of a hyperedge is $\delta_{\alpha}=\sum_{i}e_{i\alpha}$. We can think of $\delta_{\alpha}$ as the adhesiveness of $E_{\alpha}$.
Researchers have long used random walks to study hypergraphs. In Ref. Zhou et al. (2006), the authors generalized spectral clustering Shi and Malik (2000); Meilă and Shi (2000); Ng et al. (2002) from graphs to hypergraphs and gave their algorithm a random walk interpretation. It is however noted that the hypergraph random walks defined in Ref. Zhou et al. (2006) are no different than the random walks performed on expanded graphs Agarwal et al. (2006); Chitra and Raphael (2019). To take advantage of the higher order structure in hypergraphs, we need to take hyperedge degree into account when performing random walks. Since a hyperedge in a hypergraph represents an adhesive community, it is argued in Carletti et al. (2020) that a random walker roaming on a hypergraph should show preference towards hyperedges of higher degrees (stronger adhesiveness). This is not the only way to generalize random walks from graphs to hypergraphs, but it does demonstrate its advantages, as shown in Ref. Carletti et al. (2020). We also note that in Ref. Carletti et al. (2020), the authors assume that all the hypergraph nodes have the same weight, thus restricting their results only to unweighted hypergraphs. Here in this paper, we will generalize the random walks described in Ref. Carletti et al. (2020) to weighted hypergraphs, and introduce the concept of frustrated random walks on hypergraphs. We show that for heavily-weighted and scale-free hypergraphs, the definition of which will be given later, frustrated random walks are more suitable for computing node distances.
II A unified framework for calculating expected hitting times of random walks
With the abundance of research on random walks on hypergraphs, we note that the vast majority of previous work has focused on the Laplacian matrix. Spectral clustering of eigenvectors of Laplacian matrices Shi and Malik (2000); Ng et al. (2002) naturally leads to image segmentation and community detection in hypergraphs Hayashi et al. (2020); Carletti et al. (2021). Furthermore, the eigenvectors of a Laplacian matrix yield node embeddings which enable us to apply powerful machine learning algorithms to perform node classification on hypergraphs. Since the study of the Laplacian matrix yields information about the long-term stationary distribution of particles that are randomly walking on hypergraphs, we call this approach to studying random walks the stationary approach. On the other hand, we can also extract much information about hypergraphs by focusing on the diffusion process itself, and refer to this approach of studying random walks as being dynamic. In this paper, we will focus on the hitting times of random walks. Given a hypergraph $H$, we select a node $t$ as target, and perform random walks starting from any other node $s$. The hitting time for this process is defined as the number of steps a random walker needs to traverse before it hits the target $t$ for the first time. By definition, the hitting time on a any hypergraph is a random variable that depends on the hypergraph structure, the starting node $s$ and target node $t$, and can thus be denoted as $N^{(s)}_{t}$. The expectation of $N^{(s)}_{t}$ provides a natural measurement of the distance from $s$ to $t$. Note that the distance here is asymmetric, meaning that the distance from $s$ to $t$ is not guaranteed to be identical to that from $t$ to $s$. Thus, the distance here is not the distance function in the mathematical sense, and does not necessarily satisfy all the conditions, especially the symmetricity and triangle inequality, that are required of a mathematical distance function. Since we want to use the expected hitting time to measure the closeness between two persons in the real world, and we know that the real-world human-human relationships are generally asymmetric, non-equivalent, and non-reflexive, this lack of symmetry is thus considered desirable.
The dynamics of random walks on hypergraphs are fully captured by the transition matrix, and different definitions of transition matrices produce different results. We calculate hypergraph node distances to find the nearest neighbors of a node. To accomplish this, we select a node as the target, and calculate the expected hitting times starting from all other nodes. We then rank all those nodes according to their distances to the target. Intuitively, we expect that the shorter a node’s distance to the target, the closer they should be. If we represent real-world human-human interactions using a hypergraph, we expect to find a person’s close friends from among his/her nearest neighbors in the hypergraph. In this paper, we will evaluate the quality of our method’s results by checking whether a person’s close friends show up in his/her top neighbor list.
II.1 Simple random walks on hypergraphs
As already stated, the calculation of hypergraph hitting times requires specification of the random walk’s transition matrix. To capture a random walker’s preference towards a more cohesive hyperedge, we generalize the method in Ref. Carletti et al. (2020) from unweighted to weighted hypergraphs, and define transition probability from node $i$ to node $j$ as
$T^{S}_{ij,i\neq j}=\frac{\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{j\alpha}\}}{\sum_{j}\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{j\alpha}\}}$ (2)
The mechanism behind Eq. (2) is that a random walker standing on site $i$ would randomly select one of its neighbors $j$ towards which it makes a transition. A transition from $i$ to $j$ is possible only when both nodes belong to the same hyperedge. In a hypergraph that represents real-world human-human interactions, the node weight $e_{i\alpha}$ measures a member’s loyalty to the hyperedge which is interpreted as a community. Intuitively, two loyal members tend to have high frequency interactions, and two disloyal members tend not to interact with each other. The interaction frequency between a loyal member say $i$ and a disloyal member say $j$ depends on how disloyal $j$ is to the community. Thus, when calculating the transition probability from $i$ to $j$, we demand that the probability should be proportional to the lesser of the two node weights, i.e., $\min\{e_{i\alpha},e_{j\alpha}\}$, rather than their product $e_{i\alpha}e_{j\alpha}$. The proportionality coefficient $\delta_{\alpha}-e_{i\alpha}$ in front of $\min\{e_{i\alpha},e_{j\alpha}\}$ reflects the tendency of a node to stay within a more cohesive community (a hyperedge of higher degree after subtracting node $i$’s contribution). The denominator in Eq. (2) is the normalization constant.
II.2 Frustrated random walks on hypergraphs
Although simple random walk as described in Section II.1 is a good foundational algorithm for exploring hypergraph properties and can be used to generate random paths for DeepWalk, we find that the hitting times derived from this algorithm are not a good measure of node distances when the hypergraph is heavily weighted and scale-freeBarabási and Albert (1999). The fact that most real-world human-human interactions can be recast into a scale-free network compels us to develop an alternative random walk scenario which we call the frustrated random walks algorithmLi and Le (2020), an algorithm that takes into account how people interact with each other in the real world. Real world networks are mostly dominated by the Mathew effect. For example, in a social network, the more followers you have, the easier it is for you to attract new followers, a phenomenon that is called preferential attachmentBarabási and Albert (1999). The advantage of frustrated random walks over simple random walks is that it can yield node distances that are consistent with human judgment even for heavily weighted, scale-free networks. For networks that are lightly weighted or non-scale-free, results from frustrated and simple random walks tend to be similar.
In order to measure node distances on hypergraphs, we set one node (call it $t$) as the target, and perform random walks starting from another node (call it $s$). The expected hitting time of this random walk process is interpreted as the distance from $s$ to $t$. Intuitively, the shorter the distance, the closer the relationship between $s$ and $t$. We can think of this random walk process as the donation of a gift. If one person has one gift to donate, she tends to donate it to her close friend. Similarly, when a person is given more than one gift, she also tends to accept her close friend’s. With this observation, we decompose the gift-donating process into two parts: one for proposing the donation, and another for accepting it. By setting an acceptance threshold for the gift donation, we have frustrated the gift donation process, from whence arises our algorithm’s name. Under this scenario, a transition is possible only when a proposal is made and then accepted. Therefore, in frustrated random walks, the transition probability is the product of the proposal probability and the acceptance probability, which is
$T^{F}_{ij,i\neq j}$ $=\frac{\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{j\alpha}\}}{\sum_{j}\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{j\alpha}\}}$ (3)
×$\times\frac{\sum_{\beta}\Big{(}\delta_{\beta}-e_{j\beta}\Big{)}\min\{e_{j\beta},e_{i\beta}\}}{\sum_{k}\sum_{\beta}\Big{(}\delta_{\beta}-e_{j\beta}\Big{)}\min\{e_{j\beta},e_{k\beta}\}}$
Another major difference between simple random walks and frustrated random walks is that in the former case, the proposal to make a transition is always accepted and thus $T^{S}_{ii}=0,\forall i\in\mathbb{V}$, whereas in the latter case, there is a non-zero probability for the proposal to be declined, and thus $T^{F}_{ii}$ is generally non-zero. The diagonal values of $T^{F}$ matrix can be easily calculated from the observation that $\sum_{j}T^{F}_{ij}=1,\forall i,j\in\mathbb{V}$, from which we have
$T^{F}_{ii}=1-\sum_{j\neq i}T^{F}_{ij},\forall i\in\mathbb{V}.$ (4)
All of these differences combined impact significantly on the calculation of expected hitting times and ultimately the ranking of nodes with respect to a target node. We will show later, using real-world data, that frustrated random walks can better capture human-human interactions in the real world than simple random walks.
II.3 A Unified framework for calculating expected hitting times
In the previous subsections, we outlined simple random walks and frustrated random walks on hypergraphs, and formulated equations to calculate their corresponding transition probabilities. In this section, we extend the calculation to derive the expected hitting times from the transition probabilities. We show that although the transition probabilities of simple and frustrated random walks differ, the calculation of expected hitting times can be formulated in a unified framework.
For a connected hypergraph $H$, denote $N_{t}^{(s)}$ as the hitting time of a random walk process that starts from node $s$, with node $t$ as the target. We use $P(N_{t}^{(s)}=n)$ to denote the probability of reaching target node $t$ from node $s$ after exactly $n$ steps. Since the problem is trivial for the case when $s=t$, we assume the starting node $s$ differs from the target node $t$. Due to the Markov property of the random walk process, whether it be simple or frustrated, we can establish a recurrence equation for the probability, which is
$P(N_{t}^{(s)}=n)=\sum_{i_{s}\neq t}T_{s,i_{s}}P(N_{t}^{(i_{s})}=n-1),n\geq 2$ (5)
A physical interpretation of the above equation is that we decompose the random walk process into two steps. First, we make a transition from node $s$ to one of its neighbors $i_{s}$; second, due to the Markov property of the random walks, the whole process restarts all over again with $i_{s}$ as the new starting node, with the condition that now the random walker needs to reach the target with exactly $n-1$ steps since one step has already been taken. Thus, Eq. (5) is a recurrence equation for the probabilities $P(N_{t}^{(s)}=n)$. In this equation, we have imposed the condition that $n\geq 2$ because $P(N_{t}^{(j)}=0)=0$ for any node $j\neq t$, and the recursion process terminates immediately when the random walker hits the target. From the transition matrix in Eq. 2 or Eq. 3, we can readily read out the one-step transition probability $P(N_{t}^{(s)}=1)$, which is the initial condition for the recurrence relation in Eq. (5).
Eq. (5) is essentially a system of first order difference equations with constant coefficients. For simplicity of notation, we can understand $N_{t}^{(i)}$ as the $i$th component of a column vector $\boldsymbol{N}_{t}$, and $P(N_{t}^{(i)}=n)$ as the $i$th component of a column vector $\boldsymbol{X}_{n}=P(\boldsymbol{N}_{t}=n)$. With these symbols, we can rewrite Eq. (5) in an abstract form as
$\boldsymbol{X}_{n}=B\boldsymbol{X}_{n-1},n\geq 2.$ (6)
Here, $B$ is the coefficient matrix with elements depending on the transition probabilities and the target node $t$. See the appendix A for an illustrative example of how to calculate the $B$ matrix.
II.3.1 Calculation of the $B$ matrix for simple random walks
For simple random walks, $B_{ij}=T^{S}_{ij},\forall i,j\in\mathbb{V}-\{t\}$, and $B_{it}=0,\forall i\in\mathbb{V}-\{t\}$ by definition of Eq. (5). Still from Eq. (5), we have that the left subscript of $B$, which is the index of the starting node for the random walk, can never be equal to $t$. In contrast with the transition matrix $T^{S}$, matrix $B$ is not a Markov matrix because the summation of each row does not always equal one. This non-Markov property of $B$ originates from the existence of the target node $t$. From Eq. (5), we get the rule that if $t$ is not a neighbor of node $i$, then $\sum_{j}B_{ij}=1$, or else $\sum_{j}B_{ij}=1-T^{S}_{it}$. Due to this, the spectral radius of $B$ is always smaller than one.
We have established Eq. (5) to calculate the probability distribution of hitting times using any non-target node $s$ as our starting node. However, in the special case where the starting node $s$ has only one neighbor, and this very neighbor is precisely the target node $t$, then no recurrence equation is needed since we already have $P(N_{t}^{(s)}=n)=\delta_{n,1},\forall n\geq 1$. We call such a node an adherent node of target node $t$, and exclude these nodes from the set of starting nodes (For example, node 4 in Fig. 5 is an adherent to target node 3. ).
Definition 1.
In random walks, a node is called an adherent to the target if the node has the target as its only neighbor.
Since we need to exclude target node $t$ and its adherent nodes from the starting node set, the dimension of matrix $B$ is
$N_{B}=|\mathbb{V}-\{t\}|-|\{v_{i}:T^{S}_{it}=1\}|,$ (7)
where $|\mathbb{S}|$ indicates the cardinality of a set $\mathbb{S}$. Obviously, for simple random walks, the hitting time from an adherent to the target is always equal to one.
II.3.2 Calculation of the $B$ matrix for frustrated random walks
The calculation of the $B$ matrix elements for frustrated random walks is similar to that for simple random walks. Here, we still have $B_{ij}=T^{F}_{ij},\forall i,j\in\mathbb{V}-\{t\}$ and $B_{it}=0,\forall i\in\mathbb{V}-\{t\}$. The existence of target node $t$ still renders $B$ a non-Markov matrix in that $\sum_{j}B_{ij}=1$ if $i$ and $t$ do not lie simultaneously in any hyperedge, and $\sum_{j}B_{ij}=1-T^{F}_{it}$ otherwise. The only difference is that due to the imposition of the acceptance threshold for the frustrated random walks, the transition probability from an adherent to the target no longer obeys the sharp $\delta$ distribution, but a geometric distribution, i.e., $P(N_{t}^{(adherent)}=n)=(1-p)^{n-1}p$, where $p$ is the probability of $t$ accepting the transition from the adherent. Consequently, the expected hitting time from the adherent to the target is $1/p$.
II.3.3 Numerical computation of expected hitting times
Now that we have calculated the elements of the $B$ matrix, we can continue to solve Eq. (6) as
$\boldsymbol{X}_{n}=B^{n-1}\boldsymbol{X}_{1},n\geq 1.$ (8)
Here, $\boldsymbol{X}_{1}$ is the probability of hitting the target in the first step, a probability that can be conveniently read out from the transition matrix. By definition, the expected hitting times from an arbitrary node to the target node is
$\mathbb{E}\boldsymbol{N}_{t}$ $=\sum_{n=1}^{\infty}nP(\boldsymbol{N}_{t}=n):=\sum_{n=1}^{\infty}n\boldsymbol{X}_{n}$ (9)
Plugging Eq. (8) into the above definition yields an expression for the expected hitting times that depends both on the initial condition $\boldsymbol{X}_{1}$ and the coefficient matrix $B$, that is,
$\mathbb{E}\boldsymbol{N}_{t}=\sum_{n=1}^{\infty}nB^{n-1}\boldsymbol{X}_{1}$ (10)
The above equation involves an infinite summation and the powers of the $B$ matrix, and is thus difficult to compute numerically. However, we devise a convenient workaround to convert the above equation into an equivalent linear system of equations which we can efficiently solve using highly optimized linear algebra packages. To do this, we first define the probability generating function, which is
$\boldsymbol{f}(z)=\sum_{n=1}^{\infty}\boldsymbol{X}_{n}z^{n},|z|\leq 1.$ (11)
It is straightforward to show that the convergence radius of the above power series is larger than 1. More precisely, we can pinpoint the convergence radius of the series in Eq. (11) at $R=1/\rho(B)$, where $\rho(B)$ is the spectral radius of the coefficient matrix $B$. $\rho(B)$ is guaranteed to be smaller than one by the Gershgorin circle theoremMeyer (2023). With the help of Eq. (8), the probability generating function can be evaluated exactly as
$\boldsymbol{f}(z)$ $=\sum_{n=1}^{\infty}B^{n-1}\boldsymbol{X}_{1}z^{n}$ (12)
$=z(I-zB)^{-1}\boldsymbol{X}_{1},|z|\leq 1.$
Here, the summation can be evaluated in closed form because the convergence radius of the power series is larger than one. Still due to the fact that convergence radius is larger than 1, we can differentiate the infinite series in Eq. (11) term by term at $z=1$ and obtain an equivalent expression for the expected hitting times as
$\boldsymbol{f}^{\prime}(z=1)$ $=\sum_{n=1}^{\infty}nB^{n-1}\boldsymbol{X}_{1}z^{n-1}\Big{|}_{z=1}$ (13)
$=\mathbb{E}\boldsymbol{N}_{t}$
Therefore, once we have the probability generating function, we can easily calculate the expected hitting times. An example of calculating expected hitting times using this method is given in Appendix A. However, when the dimension of $B$ is huge, analytical evaluation of Eq. (11) is computationally prohibitive, which compels us to resort to numerical methods. Multiplying both sides of Eq. (12) with $I-zB$ and taking the first order derivative with respect to $z$ at $z=1$ yields
$(I-B)\boldsymbol{f}^{\prime}(z=1)=\boldsymbol{f}(z=1)$ (14)
By definition, $\boldsymbol{f}(z=1)$, which represents the sum of all probabilities, is a column vector where each element is 1. Therefore, to calculate the expected hitting times $\mathbb{E}\boldsymbol{N}_{t}=\boldsymbol{f}^{\prime}(z=1)$, we only need to solve the linear equation (14), which is straightforward to accomplish numerically. It is noteworthy that in Eq. (14), the coefficient matrix $B$ depends both on the hypergraph structure and the target node $t$, and thus target information is implicit in that equation. For real-world hypergraph data, the coefficient matrix in Eq. (14) is generally large and sparse. The best method to numerically solve a large and sparse system is the conjugate gradient methodHestenes and Stiefel (1952) or its variants, which we will use throughout our paper.
Our method for calculating hypergraph node distances can be summarized in Algorithm 1. This algorithm applies to both simple random walks and frustrated random walks. The only difference lies in the calculation of the probability transition matrix. For the sake of simplicity, we only show how to calculate the expected hitting times of simple random walks. The extension of this algorithm to frustrated random walks is straightforward.
Algorithm 1 CalculateExpectedHittingTime(H, t)
Input: hypergraph $H$
target node $t$
Return: expected hitting times $\boldsymbol{N}_{t}$
1:  Calculate transition probabilities using Eq. (2).
2:  Set node $t$ as the target.
3:  Calculate transition matrix $B$ from hypergraph structure and target node $t$, following the procedures in Section II.3.1.
4:  $d\leftarrow B.dimension$
[…]
6:  Solve $(I-B)\boldsymbol{x}=\mathbbm{1}$ using conjugate gradient method.
7:  Obtain expected hitting times $\boldsymbol{N}_{t}\leftarrow\boldsymbol{x}$
III Experimental results
In the previous sections, we have outlined two random walk algorithms, the simple random walk (SRW) and frustrated random walk (FRW), and showed that we can compute their expected hitting times in a unified framework. In this section, we use real-world hypergraph data sets to show that we can use the expected hitting times as node distances to find each node’s nearest neighbors. We also compare their results with that of DeepWalk.
III.1 Hypergraphs with ground-truth node labels
In this section, we will test our methods using two hypergraph data sets with ground truth node labels. We construct the first data set from arXiv and obtain the second data set from Ref. Chodrow et al. (2021). We note here that both hypergraphs are scale-free and lightly weighted. A hypergraph is scale-free if its node degree distribution follows a power law, and is lightly weighted if its maximum node degree is much smaller than the number of nodes in the hypergraph.
First, we create hypergraphs using articles published in arXiv under the category of physics from 2017 to 2020. Hypergraph nodes represent articles, and articles written by the same author are collected into the same hyperedge. From the hypergraph, we extract its largest connected component, which contains $52,144$ nodes (articles) and $45,188$ hyperedges (authors). The node and hyperedge degree distributions are shown in Fig. 1. In this data set, the node (i.e., article) degree represents the number of authors of the article. Here, the node degree distribution only approximately follows a power law since in recent literature most papers have more than one author. When we plot the hyperedge (i.e., author) degree distribution, a more visible power law also emerges since there is neither an upper limit nor lower limit on the number of articles that can be authored by the same researcher.
Figure 1: Log-log plot of node and hyperedge degree distribution for arXiv data set. Although the node degree distribution does not strictly follows the power law, the power law is pretty obvious for hyperedge degree distribution.
Each arXiv article has one or more subjects, which we can use as its ground truth label. Using the methods described in this paper, we can subsequently calculate hypergraph node distances, and thus find each article’s nearest neighbors. Intuitively, the shorter the distance between two articles, the more similar they should be to each other. We can quantify two articles’ similarity by calculating the Jaccard similarity between their subjects. For comparison, we find each article’s nearest neighbors using simple random walks (SRW), frustrated random walks (FRW) and DeepWalk, and calculate the mean Jaccard similarity between the target article and its top ten nearest neighbors. For one specific article, each method gives a mean Jaccard similarity score. The higher the mean Jaccard similarity, the better the method. We randomly select 12 articles, and for each method, calculate these 12 articles’ average Jaccard similarity score. The average score for each method is shown in Table 1.
Method name SRW FRW DeepWalk
Jaccard similarity 0.4092 0.4615 0.4599
Table 1: Comparison of Jaccard similarities from three methods. FRW is on par with DeepWalk, and beats SRW with a large margin.
From table 1, we can see that FRW has the highest Jaccard similarity (denoted in bold in the table). The arXiv hypergraph is approximately scale-free and lightly weighted. While FRW achieves a better result, We will further show in Section III.2 that the advantage of FRW over SRW is most obvious on heavily weighted scale-free hypergraphs.
The second data set we experiment with is the trivago-clicks data setBenson (2021). The data set has been introduced by the authors of Ref. Chodrow et al. (2021) as a collection of behavior data from users browsing trivago.com while trying to book a hotel. In this hypergraph data set, nodes represent accommodations (mostly hotels) browsed by users, and hyperedges represent a user’s browsing history in the same browsing session before checkout (i.e., before an order is placed). Each node (accommodation) is associated to a country, which can be used as the node’s label. The data set contains $172,738$ nodes and $233,202$ hyperedges. The number of distinct node labels (countries or regions) is $160$. The node degree follows a power law, yet the node degrees are generally small, ranging from $1$ to $339$, i.e., significantly smaller thatn the hypergraph node number. As a result, this hypergraph is still lightly weighted. A log-log plot of the node degree distribution is shown in Fig. 2.
Figure 2: Log-log plot of node degree distribution for trivago data set. This data set is a scale-free hypergraph.
Using SRW, FRW and DeepWalk, we can find each node’s nearest neighbors. Generally, we expect a node and its nearest neighbors to be located in the same country (having the same label). We can quantify this similarity by counting the proportion of nearest neighbors sharing the same label as that of the target node. We still randomly select 12 nodes from this hypergraph, and calculate the similarity between a target node and its top 100 nearest neighbors using the three methods. The results are shown in Table 2. On this data set, the results from the three methods are almost identical, with more than 97% nearest neighbors having the same label as that of the target node, indicating the effectiveness of all three methods in detecting nearest neighbors.
Method name SRW FRW DeepWalk
Similarity 97.08% 97.75% 97.92%
Table 2: Comparison of similarities resulting from SRW, FRW and DeepWalk methods using the trivago data set. A larger similarity indicates a better method. In this data set, all three methods give almost equally good results.
III.2 Heavily-weighted and Scale-free hypergraphs
As already stated, the node degree distribution of a scale-free hypergraph follows a power law. A hypergraph is called heavily weighted if its node degree range spans several orders of magnitude, and the maximum node degree is at least as large as the hypergraph node number. Power law distribution is ubiquitous in real world, e.g., the distributions of wealth among the world population, the number of followers of social media accounts, the citation times of scientific papers, etc. all follow a power law. According to Ref. Barabási and Albert (1999), when constructing a graph following the preferential attachment principle (which refers to attaching a newly created node to an existing node with a probability that is proportional to the node’s degree), the resultant graph will be scale-free. In a scale-free graph or hypergraph, the nodes accumulate their degrees according to the Matthew effect, meaning that the higher a node’s degree, the easier it is for its degree to get even higher. FRW is designed to capture the Matthew effect. Consequently, its advantage is most obvious for heavily-weighted and scale-free hypergraphs, as we will show in this section.
We first apply our results on a hypergraph created from Dream of the Red Chamber, a novel written by Cao Xueqin. We represent novel characters as hypergraph nodes, and each scene in the novel as a hyperedge. Characters are collected into a hyperedge if they appear in the same scene in the novel. Node degree indicates the number of times a node (character) appears in the novel. This hypergraph contains $324$ nodes and $1,524$ hyperedges. The node degree distribution follows a power law, as shown in Fig 3. The node degrees range from $1$ to $954$ which is much larger than the node number, meaning the hypergraph is heavily weighted. We will show that for such a hypergraph, the FRW method is most suitable for calculating node distances.
Figure 3: Log-log plot of node degree distribution in Dream of the Red Chamber. This figure shows the data set is a heavily-weighted, scale-free hypergraph.
Unlike the arXiv hypergraph, the novel data set does not have ground truth labels for nodes, which compels us to only compare our random walk results against results obtained from DeepWalk. The fact that DeepWalk results are consistent with human judgement further justifies this results comparison. We first run DeepWalk on the hypergraph to obtain a node ranking with respect to a target node, and then use SRW and FRW methods to get the target’s nearest neighbors. Thus, with respect to each target node, we obtain three rankings using three methods. We then calculate the rank difference between SRW/FRW results and DeepWalk results. To do so, we get the top 10 neighbors of a target node using the SRW/FRW method, find the mean DeepWalk ranking for these 10 neighbors, and finally obtain for each target node the difference between mean DeepWalk rankings and mean SRW/FRW rankings. We summarize the rank differences for several nodes in table 3. To make the results representative enough, we selected three kinds of nodes (characters) from the novel: the central characters of highest node degrees, the supporting characters of medium node degrees, and the marginal characters of low node degrees. It can be seen from the table that for this data set, FRW is much better at finding nearest neighbors of central nodes than SRW. For marginal nodes of low node degrees, the difference between FRW and SRW is negligible.
Target node Jia Baoyu Wang Xifeng Lin Daiyu Jia Zheng Jia Tanchun Jia Huan Jia Yucun Zhen Shiyin Wang Yitie
Node degree 954 554 488 279 188 94 56 22 1
SRW rank diff 99.44 86.33 41.22 74.56 53.33 43.33 0.22 0.0 0.67
FRW rank diff 7.33 0.78 4.11 4.44 4.67 16.44 0.33 0.0 2.33
Table 3: Rank difference between SRW/FRW and DeepWalk for different target nodes in Dream of the Red Chamber data set. Target nodes are ranked from high to low node degrees. For central nodes of high degrees, the differences between FRW and DeepWalk ranking results are much smaller than that for SRW, indicating that the ranking results of FRW are more consistent with DeepWalk results. Yet for marginal nodes of low degrees, the ranking results from SRW, FRW and DeepWalk are equally good.
As a second test, we create a hypergraph using the seven Harry Potter novels written by J.K. Rowling. Nodes represent novel characters, and characters that appear in the same scene are collected into the same hyperedge. This hypergraph contains $183$ nodes and $3557$ hyperedges. The maximum node degree is 4548, which is much larger than node number. Thus, this hypergraph is also heavily-weighted. In this hypergraph, the node number is too small for the power law to be obviously visible. One way to see power law in this data set is to expand the hyperedges into edges, and plot the log-log distribution of the edge weights, as shown in Fig. 4.
Figure 4: Log-log plot of edge weight distribution in Harry Potter data set. This figures shows the data set is a heavily-weighted, scale-free hypergraph.
We use SRW, FRW, and DeepWalk to find the nearest neighbors of Harry Potter, and list our results in Table 4. From the table, we can see that the results from FRW and DeepWalk are consistent with human judgment, yet SRW results are not what we expect. This corroborates our assertion that FRW is more suitable than SRW for describing real-world human relationships.
SRW FRW DeepWalk
Hokey Ron Weasley Ron Weasley
Morfin Gaunt Hermione Granger Hermione Granger
Merope Gaunt Severus Snape Albus Dumbledore
Marge Dursley Sirius Black Severus Snape
Mafalda Hopkirk Fred Weasley Ginny Weasley
Ignotus Peverell George Weasley Minerva McGonagall
Helena Ravenclaw Albus Dumbledore Voldemort
Cole Rubeus Hagrid Rubeus Hagrid
Mary Cattermole Ginny Weasley Fred Weasley
Mary Riddle Draco Malfoy Sirius Black
Table 4: Nearest neighbors of Harry Potter according to different methods. The results from FRW and DeepWalk are consistent with human judgement, and beat that from SRW.
IV Time complexity of the expected hitting time method
In Algorithm 1, we have outlined our method for calculating expected hitting times. For large hypergraphs, the most time consuming part of the algorithm is driven by finding the numerical solution of equation
$(I-B)\boldsymbol{x}=\mathbbm{1}$ (15)
For large and sparse coefficient matrix $B$, we can only solve this equation using conjugate gradient method or its variants. Conjugate gradient method involves a series of iteration cycles, and the most time consuming part of each cycle is the evaluation of $Bv$, where $B$ is the probability transition matrix and $v$ is a dense vector. Thus, the time complexity of our method as described in Algorithm 1 is $N_{iter}\times O(Bv)$, where $N_{iter}$ is the iteration number in conjugate gradient method and $O(Bv)$ is the time complexity of evaluating $Bv$. We will determine $O(Bv)$ first.
From Eq. (2) and Eq. (3), we can see that each of the summation $\sum_{j}B_{ij}v_{j}$ requires $D_{i}+1$ operations, where $D_{i}$ is the node degree of vertex $i$. Thus, the total number of operations required for computing $Bv$ is $\sum_{i}(D_{i}+1)=2E+V$, where $E$ is the edge number (the number of edges if we expand all hyperedges in the hypergraph) and $V$ is the node number. We thus have $O(Bv)=2E+V$.
To determine $N_{iter}$, we note that conjugate gradient method is guaranteed to give the exact solution to Eq. (15) after exactly $N_{B}$ iterations, where $N_{B}\lesssim V$ is the dimension of the coefficient matrix. However, for most cases, conjugate gradient method already gives results of high precision when $N_{iter}\ll N_{B}$. We can thus consider $N_{iter}$ to be a small constant number compared to $N_{B}$.
We can therefore conclude that the time complexity of our algorithm is $O(2E+V)$, with the understanding that $N_{iter}$ is a small constant number. For real-world hypergraphs, we generally have $E\propto V$. In such a case, the time complexity can be simplified to $O(V)$. Thus, the time complexity of our method is approximately linear with respect to the hypergraph size. This linear time complexity for the computation of expected hitting times of random walks is a significant advantage it has compared to the DeepWalk method.
V Comparison of running speeds
In the previous sections, we have shown that to find a node’s nearest neighbors using SRW/FRW methods, we only need to construct the transition matrix from the hypergraph structure and the target node, and solve a system of linear equations using conjugate gradient method, the time complexity of which is approximately linear. However, if we want to accomplish this using DeepWalk, we need to map all the nodes to vectors, calculate the cosine distances between the target node vector and all the other node vectors, and finally rank all the other nodes with respect to the target according to cosine distances. The node mapping could be pretty time consuming, thus rendering the DeepWalk inferior in speed compared to SRW/FRW methods.
To compare the running speeds of these methods, we test our program on three data sets (arXiv, Dream of the the Red Chamber, and Harry Potter) using a Linux system with 64 processors. When running DeepWalk, we generate random paths by walking $3200$ steps starting from each node in the hypergraph, and use word2vecMikolov et al. (2013a, b) method to map each node to a vector of size 128. To accelerate the program, the generation of random paths is fully parallelized. To obtain the running speeds for SRW and FRW methods, we randomly select some nodes as targets, and calculate the mean and standard deviation of the times spent on these nodes. The timing results are shown in Table 5.
As can be seen from the table, the running speeds of SRW and FRW are much faster than DeepWalk. SRW is even faster than FRW, yet as we have shown in the previous sections, for scale-free and heavily weighted hypergraphs, SRW cannot yield results that are on par with DeepWalk or FRW.
Data set arXiv DRC Harry Potter
DeepWalk 841 8.28 5.72
SRW $3.79\pm 0.17$ $0.012\pm 0.0002$ $0.0086\pm 0.0003$
FRW $20.12\pm 0.4$ $0.026\pm 0.0005$ $0.018\pm 0.012$
Table 5: Running times of DeepWalk, SRW and FRW on three data sets (DRC is short for Dream of the Red Chamber). The unit for all numbers is "seconds". For the task of finding one node’s nearest neighbors, SRW and FRW can beat DeepWalk in running speeds with a large margin.
VI Conclusion and outlook
In this paper, we introduced the random walk method to calculate node distances on hypergraphs, and benchmark it against DeepWalk. The random walk method includes both simple and frustrated random walks, for both of which the expected hitting times could be exactly computed in a unified framework. We used the expected hitting times of both random walk scenarios to measure node distances on hypergraphs, and tested our methods on several real-world data sets. We showed that for lightly weighted hypergraphs, simple random walk, frustrated random walk and DeepWalk give similar results, whereas for scale-free and heavily weighted hypergraphs, frustrated random walk and DeepWalk are more suitable for calculating node distances. Frustrated random walks can beat DeepWalk in running speeds when the number of target nodes is small. We finally analyzed the time complexity of our random walk method, and showed that it is approximately of linear time complexity.
Although DeepWalk can map hypergraph nodes to vectors, it is an old method that did not take advantage of the emerging graph neural networks. In the future, we want to explore the more efficient hypergraph neural networks to perform node embedding. Hopefully, this can remove our dependence on DeepWalk and accelerate our programs significantly.
Appendix A Computation of expected hitting times on an artificial hypergraph
In this section, we illustrate the procedures of calculating expected hitting times of frustrated random walks on an artificial hypergraph. Monte Carlo simulations are used to validate our numerical results. The expected hitting times of simple random walks can be calculated in a similar manner. We use the hypergraph in Fig. 5 as our example.
Figure 5: An example hypergraph, with node 3 as the target.
In this hypergraph, there are five nodes and three hyperedges. Its incidence matrix is
$e=\begin{pmatrix}1&0&0\\ 1&0&0\\ 1&1&0\\ 0&1&1\\ 0&0&1\end{pmatrix}$ (21)
We use node 3 as the target node, and calculate the expected hitting times of frustrated random walks starting from nodes {0, 1, 2, 4}. First note that, according to our definition in section II.3, node 4 is an adherent to target node 3, since node 4 has the target node as its only neighbor. From Eq. (3), we can see that the transition probability from node 4 to 3 is $p=\frac{1}{2}$. Thus, the expected hitting time from node 4 to 3 is $\mathbb{E}N_{3}^{(4)}=\frac{1}{p}=2$. Next we focus on the calculation of expected hitting times starting from nodes {0, 1, 2}.
From Eq. (3) and Eq. (5), we can establish a system of difference equations for the hitting time probabilities, which is
$\boldsymbol{X}_{n}=B\boldsymbol{X}_{n-1},n\geq 2,$ (22)
where the hitting time probability vector is
$\boldsymbol{X}_{n}=\begin{pmatrix}P(N_{3}^{(0)}=n)\\ P(N_{3}^{(1)}=n)\\ P(N_{3}^{(2)}=n)\end{pmatrix},$ (26)
and the probability transition matrix is
$B=\begin{pmatrix}11/20&1/4&1/5\\ 1/4&11/20&1/5\\ 1/5&1/5&1/2\end{pmatrix}.$ (30)
To solve Eq. (22), we need the initial condition which is $\boldsymbol{X}_{1}=\begin{pmatrix}0&0&\frac{1}{10}\end{pmatrix}^{T}$. Using the methods outlined in Section II.3.3, we can obtain the probability generating function as
$\boldsymbol{f}(z)$ $=z(I-zB)^{-1}\boldsymbol{X}_{1}$ (31)
$=\frac{1}{z(16z-65)+50}\begin{pmatrix}z^{2}\\ z^{2}\\ (5-4z)z\end{pmatrix}$ (35)
First order derivative of the above equation yields the expected hitting times, which are
$\mathbb{E}\boldsymbol{N}_{3}=\boldsymbol{f}^{\prime}(z=1)=\begin{pmatrix}35\\ 35\\ 30\end{pmatrix}$ (39)
We can similarly obtain the above results using Eq. (14), which is
$\mathbb{E}\boldsymbol{N}_{3}=(I-B)^{-1}\boldsymbol{f}(z=1)=\begin{pmatrix}35\\ 35\\ 30\end{pmatrix}$ (43)
To validate the above results, we write a Monte Carlo program to simulate the random walk process. Running the Monte Carlo simulation 100K times, we obtain the mean value of the hitting times starting from nodes {0, 1, 2, 4}. Monte Carlo simulation results and exact results are shown together in Table 6.
Expected hitting times Analytical result Monte Carlo result
$\mathbb{E}N_{3}^{(0)}$ 35 34.91879
$\mathbb{E}N_{3}^{(1)}$ 35 35.00548
$\mathbb{E}N_{3}^{(2)}$ 30 30.03275
$\mathbb{E}N_{3}^{(4)}$ 2 1.99335
Table 6: Expected hitting times of frustrated random walks on hypergraph 5 with node 3 as target, starting from nodes {0, 1, 2, 4}, using both analytical and Monte Carlo simulation methods. Analytical methods and Monte Carlo simulation yield consistent results, although analytical method can give high-precision results using much shorter computational time.
From the table, we can see that exact results are consistent with Monte Carlo simulation results, which justifies our computational method. For large hypergraphs, it is impractical to run Monte Carlo simulations because it is too time consuming. In this case, we will only use numerical methods to compute the expected hitting times.